OpenStudy (anonymous):

Which of the following correctly illustrates the conservation of mass for the reaction below? 4Na(s) + O2(g) → 2Na2O(s)

OpenStudy (taramgrant0543664):

The balanced equation is : 4 NaO2 = 2 Na2O+3 O2 molar mass NaO2 = 54.99 g/mol mass NaO2 = 54.99 g/mol x 4 mol= 220 g = mass reactant molar mass Na2O = 61.98 g/mol mass Na2O = 61.98 g/mol x 2 mol = 124 g molar mass O2 = 32 g/mol mass O2 = 32 g/mol x 3= 96 g mass products = 124 + 96 = 220 g mass reactants = mass products